OpenStudy (anonymous):

If the line passing through the points (a, 1) and (8, 7) is parallel to the line passing through the points (4, 9) and (a + 2, 1), what is the value of a?

OpenStudy (ash2326):

If the line is passing through (a,1) and (8, 7). Let its slope be m1 we know that a line passing through two points (x1, y1) and (x2, y2) is given as \[m=\frac{y2-y1}{x2-x1}\] now here the two points for first line are (a,1) and (8, 7) its slope \[m1=\frac{7-1}{8-a}\] now slope of second line m2 passing through (4,9) and (a+2, 1) is given as \[m2=\frac{1-9}{a+2-4}\] we know parallel lines have same slope so \[m1=m2\] \[\frac{7-1}{8-a}=\frac{1-9}{a+2-4}\] now we have \[\frac{6}{8-a}=\frac{-8}{a-2}\] cross multiplying we get \[6a-12=-64+8a\] bringing like terms on same side we get \[-2a=-52\] we get a \[a=26\]

OpenStudy (anonymous):

bless youu!